OpenStudy (anonymous):

GEOMETRY:the triangle that has an area of 40cm squared.,find the height N of the triangle. NUMBER THEORY: find two consecutive even integers whose product is 168. .. pls answer ,, i need it on my assignment ! :)

OpenStudy (anonymous):

is the triangle right angled?

OpenStudy (anonymous):

for number theory, the answer is 12 and 14

OpenStudy (anonymous):

assuming that the triangle is right angled, (1/2)bN=40 bN=40*2 bN=80 N=80/b

OpenStudy (anonymous):

lol I was just about to say that

OpenStudy (anonymous):

For the number theory question, even numbers are the product of 2 so let one of the two even numbers be 2x then the other one is 2(x+1) As even integers whose product is 168, (2x) 2(x+1)=168 x(x+1)=42 x^2 +x -42=0 (x+7)(x-6)=0 x=-7 or 6 For x=-7, the two numbers are -14 or -12 For x=6, the two numbers are 12 or 14

OpenStudy (anonymous):

@jesamaeadvincula

OpenStudy (anonymous):

can you see this page, @jesamaeadvincula